4 ,1 ,-2
Which term of the sequence is equal to -143

Answer by jim_thompson5910(35256)

First term: a1 = 4

Common Difference: d = -3

nth term:

an = a1 + d(n-1)

an = 4 + (-3)(n-1)

an = 4 - 3n + 3

an = -3n + 7

Plug in an = -143 and solve for n

an = -3n + 7

-143 = -3n + 7

-143-7 = -3n

-150 = -3n

-150/(-3) = n

50 = n

n = 50

So it's the 50th term.